A golf ball is driven horizontally from an elevated tee with a velocity of 27 m/s. It strikes the fairway 2.5 sec later. How far has it fallen horizontally?


1
Expert's answer
2021-10-18T03:44:01-0400

"y = ut+ \frac12at²\\
y = 0(t) + \frac12(9.8)+(2.5)²\\
y = 0+ 4.9(6.25)\\
y = 30.625m"


"\therefore" The vertical distance is 30.625m